### v1

@@ 1 Introduction
Persistent currents are frictionless flows of quantum matter occurring in annular geometries pierced by an effective magnetic field. The existence of persistent currents was discovered in superconducting circuits at the turn of the 20th century [1, 2, 3, 4, 5]. Such a notion represented a milestone for the development of the concept of superfluidity [6, 7, 8, 9]. Rather than zero viscosity or zero resistance, the persistent current arises from the system being coherent beyond of a microscopic length scale  [10, 11]. Indeed, although smaller than the superconducting/superfluid counterparts, dissipationless electric currents has been observed also in normal metallic settings as the electrons’ coherence length gets comparable with the characteristic size of the ring [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. As a function of the threaded magnetic flux, specific Aharonov-Bohm oscillations can emerge [22, 23, 17, 12, 14, 13]. Persistent currents can only be observed in the quantum regime at very low temperatures where decoherence effects coming for example from disorder and thermal fluctuations are negligible - see f.i.[24, 25, 26, 27, 28]. As a result, detecting persistent currents in normal metals has turned out to be quite challenging [29, 24, 30, 31].
@@
Soon after the first experiments on bosonic ultracold atoms [32, 33, 34, 35], attention was devoted to their superfluid features [36, 37, 38, 39, 40, 41]. Despite being charge-neutral in nature, ultracold atoms can be set in motion through the application of artificial gauge fields that can be realized with different approaches [42, 43], ranging from stirring that exploits the equivalence between the Lorentz and Coriolis forces [44, 45, 46, 47] to suitable Raman transitions [48, 49, 50, 51] and phase imprinting techniques employing nowadays commercially available devices for light sculpting such as Spatial Light Modulators (SLM) or Digital Mirror Devices (DMD)  [52, 53, 54, 55, 56, 57, 58]. Ultracold atoms provide a platform to realize persistent currents with very specific capabilities that are hardly, if not impossible, to be accessed by standard schemes, as solid-state physics implementations. To start with, ultracold atoms have paradigmatically robust coherence properties and control of the physical conditions [59, 60, 61]. Sophisticated high-resolution image systems allow to monitor the system with a remarkably high precision - see f.i. [62, 63, 64, 65, 66, 67, 68]. Dynamics of ultracold matter can be studied to an unprecedented degree since the present experimental know-how allow to change the system’s physical conditions during the course of the experiment ‘on the fly’ [55, 69, 70]. Furthermore, the wide range of atomic species that can be used in cold atom experiments allows to study persistent current with a large variety of different quantum fluids, as $N$-component bosons/fermions or mixtures of thereof  [71, 72, 73, 74, 75, 76, 77, 78]. their quantization properties have been addressed directly in a number of experiments. Specifically, the earliest experimental studies of matter-wave currents in ultracold systems involved weakly interacting bosonic atoms [51, 49, 50, 46, 79, 80, 81], but work in this area recently expanded to include fermionic-pair superfluids with tunable interactions [47, 53, 82, 83]. Persistent currents of $N$-component gases [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], i.e. atoms with $N$ internal degrees of freedom, and of Bose-Bose [95, 96, 97] or Bose-Fermi mixtures [98] have been investigated theoretically. Close networks of Rydberg atoms provide a conceptual extension of the field and have attracted recent attention [64, 99, 100, 69]. In such proposals, rather than of matter, a controlled current can occur in terms of atomic excitations [101, 102, 103, 104, 105, 106, 107, 108].
@@
Finally, persistent currents of neutral matter-waves convey a technological significance. In particular they provide a key ingredient of Atomtronics, the quantum technology of ultracold atoms propagating in magneto-optical guides [126, 127, 128, 129, 130]. Indeed, ring-shaped degenerate gases interrupted by few localized weak links defines the atomic analogs of Superconducting Quantum Intereference Devices (SQUIDs) of quantum electronics [131]. The ultracold atoms conterpart of SQUIDs have been experimentally realized [132, 122, 123, 133, 134]. In virtue of the notoriously mild decoherence properties of such platforms, central questions for quantum science such as the macroscopic quantum tunneling/coherence [135], have been studied intensively [136, 137, 138, 139, 140, 141, 122, 142, 123, 114]. At the same time, atomic SQUIDs are very promising in quantum sensing [143] as compact Sagnac interferometers [144, 145, 146, 147] or gyroscopes [148, 149, 150, 151, 152, 153, 147, 154].
The review is organized as follows. In Sec. 2, we discuss how persistent currents can arise in mesoscopic cold atom rings and address few relevant features on the concept of superfluidity and phase coherence. Sec. 3 introduces the theoretical framework used to understand persistent currents in bosonic and fermionic systems. The ensuing sections (Secs. 4 and 5) address persistent currents in bosonic and fermionic systems, respectively. In Sec. 6, we discuss persistent currents can set the basis for relevant atomtronics applications in quantum technologies as rotation sensing, persistent current-based qubits and controlled manipulations of Rydberg atoms. Finally, we conclude our review in Sec. 7 by presenting few remarks and expounding on the future prospects for the field.
@@ 2 Concept of persistent currents in cold atoms
Persistent currents are a manifestation of quantum coherence. On the operational side, we note that such phenomenon is achieved by promoting the ground state of the system at rest to an excited state in which particles circulate in the toroidal track. As such, persistent currents should be understood as metastable states with a very long life time [41, 9]. This important concept relies on having a finite critical velocity in the system. In particular, for the ideal Bose gas that has a zero critical velocity, the dissipation-free current decays once the synthetic gauge field is removed. Indeed, the persistent current can be expressed as a thermodynamic equilibrium function only in a frame in which the synthetic field is at rest –see Sec. 2.1. For persistent currents achieved by stirring, for example, the latter is provided by the frame rotating with the stirring laser potential –see Sec .2.3. Cold atoms settings have allowed to highlight the stability features of persistent currents with remarkable insights –see Sec. 4.4.
In this section, we provide the main ingredients of the persistent current phenomena in cold atoms. Relations with superfluidity and phase coherence give important insights on the concept of persistent currents. Being that the current is comprised of electrically neutral particles, specific schemes for applying effective magnetic fields have been experimentally realized. Persistent currents are nowadays accessible in several experimental laboratories through suitable interference protocols.
@@ 2.1 Persistent currents in ultracold atoms
Here, we illustrate how subjecting cold atoms confined in mesoscopic rings to a synthetic gauge field triggers dissipationless quantized matter-wave currents. Our approach follows that of electrons in clean metallic rings carried out in [155].
The persistent current in mesoscopic metallic rings is generated by applying a static uniform magnetic field, f.i. $\mathbf{B}=B\mathbf{z}$ [12, 13, 14]. In the presence of a magnetic field, the kinetic momentum of the electrons $p$ is modified due to its coupling with the magnetic vector potential $\mathbf{A}$ through their charge $q$, with $\mathbf{B}=\nabla\times\mathbf{A}$. Employing the Lagrangian for a single particle in an electromagnetic field, the canonical momentum takes the form $\textbf{P}=\textbf{p}+q\mathbf{A}$. The corresponding Hamiltonian expressed in terms of canonical momenta and coordinates reads
@@
where $\Phi(\mathbf{r})$ is a free gauge that can be taken to be zero [156]. Naturally, the kinetic momentum of ultracold neutral atoms is not affected by external magnetic fields. However, it is possible to engineer synthetic gauge fields through various means [42, 43] such that the behavior of charged particles subjected to Lorentz forces in presence of a magnetic field is emulated in a cold atom system, leading to a Hamiltonian bearing a similar form to Eq. (1)
@@
The specific form of $\chi$ will depend on the scheme utilized to generate the synthetic gauge fields, some of which will be discussed in Sec. 2.3.
To understand the implications of the artificial gauge field, we first revisit the time-independent Schrödinger equation for a single quantum particle having a mass $m$ residing in a one-dimensional ring of radius $R$
$-\frac{\hbar^{2}}{2mR^{2}}\frac{\partial^{2}}{\partial\varphi^{2}}\psi_{n}(\varphi)=\epsilon_{n}\psi_{n}(\varphi),$ (3)
where $\varphi$ corresponds to the azimuthal polar coordinate. The wavefunction $\psi_{n}=\frac{1}{\sqrt{2\pi R}}e^{\imath n\varphi}$ are plane waves having an associated angular momentum $n\hbar$ and corresponding energy $\epsilon_{n}=\frac{\hbar^{2}}{2mR^{2}}n^{2}$, which can be readily obtained by imposing periodic boundary conditions on the wavefunction, $\psi(\varphi)=\psi(\varphi+2\pi)$.
@@
$\frac{1}{2m}\bigg{(}\frac{-\imath\hbar}{R}\frac{\partial}{\partial\varphi}-\frac{\chi\phi}{2\pi R}\bigg{)}^{2}\psi_{n}(\varphi)=\epsilon_{n}\psi_{n}(\varphi),$ (4)
on expressing the vector potential, that we defined as $\mathbf{A}=\frac{1}{2}\mathbf{B}_{a}\times\mathbf{r}$, in terms of the flux $\phi$ threading the ring, with $\phi=\pi r^{2}B_{a}$. The field is chosen such that it is null at the ring’s radius, indicated by $R_{c}$ in Fig. 1. We remark that even if the field extends beyond this region $R>R_{c}$, the effects will only depend on the flux piercing the ring. Whilst the eigenfunctions are still plane waves as before, i.e. independent of the flux, the corresponding energies acquire the following form
@@
where we defined the elementary flux quantum as $\phi_{0}=\frac{h}{\chi}$. Plotting the spectrum against the flux (see Fig. 2(a)), one finds that the energy of a given eigenstate is a parabola due to the quadratic nature of the energy, implying in turn the symmetry $\phi\leftrightarrow-\phi$. Despite this, the energy spectrum is piece-wise parabolic, periodic with the flux having a period $\phi_{0}$ akin to particles in a periodic potential as per the Bloch theorem. This occurs on account of crossings between levels of different $n$ at $\phi=\frac{j}{2}\phi_{0}$ for integer $j$: as the flux piercing the ring increases, there is a shift in the angular momentum of the circulating particles counteracting it, causing these intersections.
@@
Before moving forward, we point out that instead of accounting for the flux dependence of the system through the Hamiltonian, one could have included it in an alternative way. Performing a unitary transformation, with a suitably chosen gauge, enables us to model the system with a field-free Hamiltonian and incorporate the influence of the flux in the periodic boundary conditions such that the boundaries become twisted. To carry out this operation, we modify the earlier definition of the vector potential to $A=\nabla\xi$, satisfying our constraint of having a zero magnetic field at the ring’s radius. For a one-dimensional ring, as is the case here, one can choose $\xi=\frac{\phi}{2\pi}\varphi$, that when employed through the unitary operator $\mathcal{U}=e^{-\imath(\chi/\hbar)\xi}$ on the Hamiltonian
$\mathcal{U}\bigg{(}-\frac{\imath\hbar}{R}\frac{\partial}{\partial\varphi}\bigg{)}\mathcal{U}^{-1}=-\frac{\imath\hbar}{R}\frac{\partial}{\partial\varphi}+\frac{\chi\phi}{2\pi R},$ (6)
nullifies the vector potential reverting back to the Hamiltonian in its field-free form [155]. On the other hand, an overall phase becomes encoded in the wavefunction $\psi^{\prime}=\psi\exp(-\frac{\imath\chi}{\hbar}\xi)$. Generally, this acquired factor bears no impact when calculating physical observables. Nevertheless, a coherent system where particles can traverse along two possible trajectories winding around a flux produces an interference pattern that is shifted by the induced phase. Interestingly enough, this occurs despite the particles moving in a region devoid of the artificial magnetic field as it solely originates from the interaction with the gauge field. Such a phenomenon is due to the celebrated Aharonov-Bohm effect, with the considered scenario being an instance of this phenomenon in a closed loop [22, 157]. The so-called Aharonov-Bohm phase picked up by the atoms is proportional to the gauge field and so by definition it turns out to be
$-\imath\frac{\chi}{\hbar}\Delta\xi=\oint\limits_{2\pi R}A\cdot\mathrm{d}l=-2\pi\imath\frac{\phi}{\phi_{0}}.$ (7)
As a result, at the boundary we have $\psi^{\prime}(\varphi+2\pi)=e^{-2\pi\imath\frac{\phi}{\phi_{0}}}\psi^{\prime}(\varphi)$ with this new constraint being the twisted boundary condition mentioned previously. Through this modified boundary it becomes saliently clear that the spectrum, which is naturally conserved under the unitary transformation, is periodic in the flux $\epsilon(n^{\prime})=\epsilon\big{(}n^{\prime}+\frac{\phi}{\phi_{0}}\big{)}$ with a period dictated by $\phi_{0}$.
Our starting point in understanding how a persistent current emerges from the Aharonov-Bohm effect is the typical current definition: the net rate of flow of atoms past a given point $I=N_{p}/\Delta t$ [^1] In the case of metals, the current needs to be multiplied by the electric charge $e$.. Consider the average current for a single particle taken over one period, $\Delta t=L/v$, with $v=p/m$ being the particle’s velocity around the ring and $L=2\pi R$. Utilising the previously established definition of the kinetic momentum
$v_{n}\psi_{n}(\varphi)=\frac{1}{m}(P-\chi\mathbf{A})\psi_{n}(\varphi)\implies v_{n}=\frac{h}{mL}\bigg{(}n-\frac{\phi}{\phi_{0}}\bigg{)},$ (8)
gives us the current associated with each eigenstate
$I_{n}(\phi)=\frac{h}{mL^{2}}\bigg{(}n-\frac{\phi}{\phi_{0}}\bigg{)}.$ (9)
Going a step further, the current can be expressed in terms of the energy via the thermodynamic relation $I_{n}(\phi)=-\frac{1}{h}\partial\epsilon_{n}/\partial\tilde{\phi}$, where $\tilde{\phi}=\phi/\phi_{0}$ [^2] We remark that $\chi$ is incorporated into the current to account for the physical nature of the system. For example, the charge current in metals is obtained by setting $\chi=e$ and for cold atoms matter-wave currents generated through rotations, as will be discussed in Sec. 2.3, $\chi=mR^{2}$ to account for the moment of inertia of a rotating body [44, 41]. Dividing this by $\chi$ gives the typical expression for the probability current (density).. One could illustrate this by going about it in a different way. Firstly, we return to Eq. (4), multiply both sides from the left by $\psi^{*}_{n}$ and subsequently integrate along the perimeter of the ring leaving us with
$\epsilon_{n}(\phi)=-\frac{\hbar^{2}}{2m}\int\limits_{0}^{2\pi}\psi_{n}^{*}(\varphi)K^{2}\psi_{n}(\varphi)R\hskip 2.84526pt\mathrm{d}\varphi,$ (10)
wherein we defined $K=\frac{1}{R}\frac{\partial}{\partial\varphi}-\frac{\imath\chi}{\hbar}\mathbf{A}$ for the sake of convenience. Performing a derivation with the flux followed by integration by parts [155] gives that
$\frac{\partial\epsilon_{n}}{\partial\phi}=\frac{1}{2\pi R}\frac{\imath\chi\hbar}{2m}\int\limits_{0}^{2\pi}[\psi_{n}^{*}(\varphi)K\psi_{n}(\varphi)-\psi_{n}(\varphi)K^{*}\psi^{*}_{n}(\varphi)]R\hskip 2.84526pt\mathrm{d}\varphi.$ (11)
The spatially averaged particle current is formulated as $I_{n}(\phi)=\frac{1}{L}\int^{L}_{0}j_{n}(x,\phi)\mathrm{d}x$ with $j_{n}(x,\phi)$ being the definition of the current density in quantum mechanics for particles in an artificial gauge field
$j_{n}(x,\phi)=\frac{\chi}{2m}\bigg{[}(\Psi_{n}^{*}p\Psi_{n}-\Psi_{n}p\Psi_{n}^{*})-2\chi\mathbf{A}|\Psi^{2}|\bigg{]}.$ (12)
Putting everything together, we have that $-\frac{\partial\epsilon_{n}}{\partial\phi}=\frac{1}{2\pi}\int_{0}^{2\pi}j_{n}(\varphi,\phi)\mathrm{d}\varphi\equiv I_{n}(\phi)$, which is the same expression that was found previously. In principle, to calculate the persistent current one needs to look at the contribution of all the eigenstates, and then sum them up after introducing a suitable weighting factor that reflects the thermal population of the states, i.e. the Fermi-Dirac distribution $\mathcal{F}(\epsilon_{n},T)$, such that the total current reads $I(\phi)=\sum_{n}I_{n}\mathcal{F}(\epsilon_{n},T)$. A straightforward calculation leads us to
$I(\phi)=-\frac{1}{h}\frac{\partial F(\phi)}{\partial\tilde{\phi}}\;,$ (13)
where $F$ denotes the Helmholtz free energy of the canonical system [16]. Being a derivative of the thermodynamic potential, we emphasize that the persistent current in the frame in which the synthetic field is at rest is an equilibrium thermodynamic physical quantity. Naturally, at zero temperature, only the ground-state level of the system is occupied such that expression for the persistent current reads
$I_{0}(\phi)=-\frac{1}{h}\frac{\partial E_{0}(\phi)}{\partial\tilde{\phi}},$ (14)
@@
where $E_{0}$ denotes the ground-state energy of the system and $I_{0}$ is the corresponding circulating current [^3] For cold atom systems, the main contribution to the current is from the ground-state. Going forward, the ground-state current will be denoted by $I$.. Although the expression for the persistent current obtained in (14) is for the single-particle case, it also holds true for a many-body system. Naturally in the absence of interactions, this is a straightforward generalization by summing up the different contributions of the single-particle currents over the particle distribution $\{n\}$ that for a system of $N_{p}$ particles reads $I(\phi)=\sum_{\{n\}}\frac{\hbar}{mR^{2}}\big{(}n-\frac{\phi}{\phi_{0}}\big{)}$. For a system consisting of $N_{p}$ particles, we define =$\ell=\sum_{\{n\}}n/N_{p}$ as the angular momentum per particle or as is most commonly referred to in experiments the winding number. In the case of interacting systems, the expression for the persistent current is the same as the free particle regime, which albeit less clear, follows from the fact that the flux dependence of the system is encoded solely in the kinetic term (easily seen by using the Peierls substitution). This reflects the notion that through a linear combination of Slater determinants, one can obtain any circulating current state [155].

### v3

@@ 1 Introduction
Persistent currents are permanent flows of quantum matter occurring in annular geometries pierced by an effective magnetic field. The existence of persistent currents was discovered in superconducting circuits at the turn of the 20th century [1, 2, 3, 4, 5]. Such a notion represented a milestone for the development of the concept of superfluidity [6, 7, 8, 9]. Rather than zero viscosity or zero resistance, the persistent current arises from the system being coherent beyond of a microscopic length scale  [10, 11]. Indeed, although smaller than the superconducting/superfluid counterparts, dissipationless electric currents has been observed also in normal metallic settings as the electrons’ coherence length gets comparable with the characteristic size of the ring [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. As a function of the threaded magnetic flux, specific Aharonov-Bohm oscillations can emerge [23, 24, 17, 12, 14, 13]. Persistent currents can only be observed in the quantum regime at very low temperatures where decoherence effects coming for example from disorder and thermal fluctuations are negligible - see for instance [25, 26, 27, 28, 29, 30, 31, 32]. As a result, detecting persistent currents in normal metals has turned out to be quite challenging [33, 25, 34, 35].
@@
Soon after the first experiments on bosonic ultracold atoms [36, 37, 38, 39], attention was devoted to their superfluid features [40, 41, 42, 43, 44, 45]. Despite being charge-neutral in nature, ultracold atoms can be set in motion through the application of artificial gauge fields that can be realized with different approaches [46, 47], ranging from stirring that exploits the equivalence between the Lorentz and Coriolis forces [48, 49, 50, 51] to suitable Raman transitions [52, 53, 54, 55] and phase imprinting techniques employing nowadays commercially available devices for light sculpting such as Spatial Light Modulators (SLM) or Digital Micromirror Devices (DMD)  [56, 57, 58, 59, 60, 61, 62]. Ultracold atoms provide a platform to realize persistent currents with very specific capabilities that are hardly, if not impossible, to be accessed by standard schemes, as solid-state physics implementations. To start with, ultracold atoms have paradigmatically robust coherence properties and control of the physical conditions [63, 64, 65]. Sophisticated high-resolution image systems allow to monitor the system with a remarkably high precision - see for example [66, 67, 68, 69, 70, 71, 57]. Dynamics of ultracold matter can be studied to an unprecedented degree since the present experimental know-how allows to change the system’s physical conditions during the course of the experiment ‘on the fly’ [59, 72, 73]. Furthermore, the wide range of atomic species that can be used in cold atom experiments allows to study persistent current with a large variety of different quantum fluids, as $N$-component bosons/fermions or mixtures thereof [74, 75, 76, 77, 78, 79, 80, 81]. Indeed, atomic persistent currents together with their quantization properties have been addressed directly in a number of experiments. Specifically, the earliest experimental studies of matter-wave currents in ultracold systems involved weakly interacting bosonic atoms [55, 53, 54, 50, 82, 83, 84], but work in this area recently expanded to include fermionic-pair superfluids with tunable interactions [51, 57, 85, 86]. Persistent currents of $N$-component gases [87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100], i.e. atoms with $N$ internal degrees of freedom, and of Bose-Bose [101, 102, 103] or Bose-Fermi mixtures [104] have been investigated theoretically. Close networks of Rydberg atoms provide a conceptual extension of the field and have attracted recent attention [68, 105, 106, 72]. In such proposals, rather than of matter, a controlled current can occur in terms of atomic excitations [107, 108, 109, 110, 111, 112, 113, 114].
@@
Finally, persistent currents of neutral matter-waves convey a technological significance. In particular they provide a key ingredient of Atomtronics, the quantum technology of ultracold atoms propagating in magneto-optical guides [132, 133, 134, 135, 136]. Indeed, ring-shaped degenerate gases interrupted by few localized weak links define the atomic analogs of Superconducting Quantum Intereference Devices (SQUIDs) of quantum electronics [137]. The ultracold atoms counterpart of SQUIDs have been experimentally realized [138, 128, 129, 139, 140]. In virtue of the notoriously mild decoherence properties of such platforms, central questions for quantum science such as the macroscopic quantum tunneling/coherence [141], have been studied intensively [142, 143, 144, 145, 146, 147, 128, 148, 129, 120]. At the same time, atomic SQUIDs are very promising in quantum sensing [149] as compact Sagnac interferometers [150, 151, 152, 153] or gyroscopes [154, 155, 156, 157, 158, 159, 153, 160].
The review is organized as follows. In Sec. 2, we discuss how persistent currents can arise in mesoscopic cold atom rings and address few relevant features on the concept of superfluidity and phase coherence. Sec. 3 introduces the theoretical framework used to understand persistent currents in bosonic and fermionic systems. The ensuing sections (Secs. 4 and 5) address persistent currents in bosonic and fermionic systems, respectively. In Sec. 6, we discuss how persistent currents can set the basis for relevant atomtronics applications in quantum technologies such as rotation sensing, persistent current-based qubits and controlled manipulations of Rydberg atoms. Finally, we conclude our review in Sec. 7 by presenting few remarks and expounding on the future prospects for the field.
@@ 2 Concept of persistent currents in cold atoms
Persistent currents are a manifestation of quantum coherence. On the operational side, we note that such phenomenon is achieved by promoting the ground state of the system at rest to an excited state in which particles circulate in the toroidal track. As such, persistent currents should be understood as metastable states with a very long life time [45, 9]. This important concept relies on having a finite critical velocity in the system. In particular, for the ideal Bose gas that has a zero critical velocity, the dissipation-free current decays once the synthetic gauge field is removed. Indeed, the persistent current can be expressed as a thermodynamic equilibrium function only in a frame in which the synthetic field is at rest - see Sec. 2.1. For persistent currents achieved by stirring, for example, the latter is provided by the frame rotating with the stirring laser potential - see Sec. 2.3. Cold atoms settings have allowed to highlight the stability features of persistent currents with remarkable insights- see Sec. 4.4.
In this section, we provide the main ingredients of the persistent current phenomena in cold atoms. Relations with superfluidity and phase coherence give important insights on the concept of persistent currents. Since the current is comprised of electrically neutral particles, specific schemes for applying effective magnetic fields have been experimentally realized. Persistent currents are nowadays accessible in several experimental laboratories through suitable interference protocols.
@@ 2.1 Persistent currents in ultracold atoms
The purpose of this section is to demonstrate that the phenomenon of persistent currents arises as a manifestation of the phase of the wavefunction. Here, we illustrate how subjecting cold atoms confined in mesoscopic rings to a synthetic gauge field triggers quantized matter-wave currents. Our approach follows the logic carried out in [161] adapting it to neutral atoms.
In the presence of a magnetic field $\mathbf{B}$, the canonical momentum of the electrons is modified due to its coupling with the magnetic vector potential $\mathbf{A}$ through their charge $q$, with $\mathbf{B}=\nabla\times\mathbf{A}$. Employing the Lagrangian formalism the canonical momentum can be demonstrated to be $\textbf{P}=\textbf{p}+q\mathbf{A}$, where $\textbf{p}=m\textbf{v}$ is the kinetic momentum. The corresponding Hamiltonian expressed in terms of canonical momenta and coordinates reads
@@
where $\mathbf{P}=-\imath\hbar\nabla$ and $\Phi(\mathbf{r})$ is the scalar potential that can be taken to be identically zero in the radiative gauge [162]. Naturally, the kinetic momentum of ultracold neutral atoms is not affected by external magnetic fields. However, it is possible to engineer synthetic gauge fields through various means [163, 46, 47] such that the behavior of charged particles subjected to Lorentz forces in presence of a magnetic field is emulated in a cold atom system, leading to a Hamiltonian bearing a similar form to Eq. (1)
@@
The specific form of $\chi$ will depend on the scheme utilized to generate the synthetic gauge field, some of which will be discussed in Sec. 2.3. To understand the implications of the artificial gauge field, we first revisit the time-independent Schrödinger equation for a single quantum particle having a mass $m$ residing in a one-dimensional ring of radius $R$. In this geometry, $\nabla=\frac{1}{R}\frac{\partial}{\partial\varphi}$ where $\varphi$ corresponds to the azimuthal polar coordinate.
@@
$\frac{1}{2m}\bigg{(}\frac{-\imath\hbar}{R}\frac{\partial}{\partial\varphi}-\chi A_{\varphi}\bigg{)}^{2}\psi_{n}(\varphi)=\epsilon_{n}\psi_{n}(\varphi),$ (3)
where we expressed the vector potential as $\mathbf{A}=\frac{1}{2}\mathbf{B}_{a}\times\mathbf{r}=A_{\varphi}\mathbf{\hat{\varphi}}$, where $A_{\varphi}$ is determined by imposing the circular integral $R\oint A_{\varphi}\mathrm{d}\varphi=\phi$, with $\phi$ being the magnetic flux threading the ring: $\phi=\pi R^{2}B_{a}$. The field is chosen such that it is null at the ring’s radius, as indicated in Fig. 1. We remark that even if the field extends beyond its parameter, the effects will only depend on the flux piercing the ring. The eigenfunctions $\psi_{n}=\frac{1}{\sqrt{2\pi R}}e^{\imath n\varphi}$ satisfying Eq. (3) are plane waves where $n$ is an integer fixed by imposing periodic boundary conditions on the wavefunction, $\psi(\varphi)=\psi(\varphi+2\pi)$, and associated with an angular momentum $n\hbar$. The corresponding energies are:
@@
where we defined the elementary flux quantum as $\phi_{0}=\frac{h}{\chi}$. Plotting the spectrum against the flux (see Fig. 2(a)), one finds that $\epsilon_{n}$ is is piece-wise parabolic, with a symmetry $\phi\leftrightarrow-\phi$. Therefore, $\epsilon_{n}$ turns out to be a periodic function of $\phi$ with period $\phi_{0}$. Specifically, there are crossings between levels of different $n$ at $\phi=\frac{j}{2}\phi_{0}$ or $\phi=j\phi_{0}$ for integer $j$ depending on the particle number parity: as the flux piercing the ring increases, there is a shift in the angular momentum of the circulating particles counteracting it, causing these intersections.
@@
In what follows, we make the connection with the Aharonov-Bohm effect, which generally results from a specific quantum interference experienced by particles propagating in a solenoidal magnetic field [23, 164] (for the experimental demonstration of the effect see [165, 166]). To this end, we perform a gauge transformation that, as we shall see, removes the vector potential from the Hamiltonian and incorporates a phase factor in the wavefunction. This is implemented by rotating the Hamiltonian in Eq. (3) using $\mathcal{U}K\mathcal{U}^{-1}\mathcal{U}K\mathcal{U}^{-1}\mathcal{U}\psi=-\frac{2m}{\hbar^{2}}\epsilon_{n}\;\mathcal{U}\psi$ where $K=\left(\frac{1}{R}\frac{\partial}{\partial\varphi}-\frac{\imath\chi}{\hbar}A_{\varphi}\right)$ and $\mathcal{U}=e^{-\frac{\imath}{\hbar}\chi R\int A\mathrm{d}\varphi}$:
[…]
Correspondingly, the wavefunction acquires a phase factor $\psi\rightarrow\psi^{\prime}=e^{-\frac{i}{\hbar}\chi R\int A\mathrm{d}\varphi}\psi$, which coincides with the Aharonov-Bohm phase. Additionally, such a phase also modifies the boundary conditions of the system
$\psi^{\prime}(\varphi+2\pi)=e^{-\frac{i}{\hbar}\chi R\oint A\mathrm{d}\varphi}\psi^{\prime}(\varphi)\,.$ (6)
The response of the system to the change in the boundary conditions was studied by Kohn [167] and further discussed by Thouless and co-workers [168, 169]. We note that making a correspondence between the crystal momenta and the effective magnetic flux $\phi$, the expression of the wavefunction Eq. (6) makes the basis for the formal analogy between particles in ring pierced by a magnetic field and the Bloch theory describing particles in periodic potentials [170]. In this sense, as originally noted by Büttiker, Imry and Landauer [12, 21], and discussed by Leggett [171], $\epsilon(\phi)$ provides the ‘band structure’ for the problem – See also Sec. 3.3
To obtain the current $I(\phi)=-\frac{1}{L}\int^{L}_{0}j_{n}(x,\phi)\mathrm{d}x$ we refer to the probability current density:
$j_{n}(x,\phi)=\frac{\imath\hbar}{2m}\bigg{[}[\psi_{n}^{*}(\varphi)K\psi_{n}(\varphi)-\psi_{n}(\varphi)K^{*}\psi^{*}_{n}(\varphi)\bigg{]}=-\frac{h}{mL^{2}}\bigg{(}n-\frac{\phi}{\phi_{0}}\bigg{)},$ (7)
which results in:
$I(\phi)=\frac{h}{mL^{2}}\bigg{(}n-\frac{\phi}{\phi_{0}}\bigg{)}.$ (8)
We note that Eq. (8) can also be obtained through the derivative of the energy Eq. (4):
$I(\phi)=-\frac{1}{h}\frac{\partial\epsilon_{n}(\phi)}{\partial\tilde{\phi}},$ (9)
where $\tilde{\phi}=\phi/\phi_{0}$. A relation similar to Eq. (9) holds for finite temperatures. Specifically, at finite temperatures, the current emerges from the occupation of the eigenstates given by their thermal distribution and thus can be obtained from the the Helmholtz free energy $F$ as $I_{th}(\phi)=-\frac{1}{h}\frac{\partial F(\phi)}{\partial\tilde{\phi}}$. Clearly, at zero temperature, only the ground-state level of the system is occupied such that the expression for the current reads
$I(\phi)=-\frac{1}{h}\frac{\partial E_{0}(\phi)}{\partial\tilde{\phi}},$ (10)
where $E_{0}$ denotes the ground-state energy of the system and $I$ is the corresponding circulating current.
In summary, persistent currents arise from the coupling between the effective magnetic field and the phase of the wavefunction. This coupling leads to a sizable dependence of the system’s energy on $\phi$, for example quantified by the $\phi$-bandwidth. Moreover, in the frame in which the synthetic field is at rest, the persistent current is an equilibrium quantity, thus, it never decays as long as $\phi$ is present. In real systems, it should be clear that a persistent current can emerge only if the phase can be preserved during the particles’ motion on a characteristic length scale [168, 172, 173]. Such a property, defining the notion of phase coherence, will be further elaborated in the next section. In particular, here we note that for ultracold atoms the persistent current can be started also by a suitable engineering of the phase of the ‘macroscopic’ wavefunction (see also Sec. 2.3).
@@
Although the theory discussed above refers to the single-particle case, the significance and most of general aspects of the persistent current we worked out can be applied to interacting many-body systems. For a system of $N_{p}$ particles $I(\phi)=\sum_{\{n\}}\frac{h}{mL^{2}}\big{(}n-\frac{\phi}{\phi_{0}}\big{)}$, which is obtained by adding the different contributions of the single-particle currents over the particle distribution $\{n\}$. The corresponding angular momentum per particle is then defined as =$\ell=\sum_{\{n\}}n/N_{p}$, which is also commonly referred to in experiments as the winding number. Additionally, we note that the flux dependence of the system is encoded solely in the kinetic term. This reflects the notion that through a linear combination of Slater determinants, one can obtain any circulating current state [161].